OpenStudy (zeesbrat3):

A particle moves along the x-axis with position function s(t) = xe^x. How many times in the interval [−5, 5] is the velocity equal to 0?

OpenStudy (zeesbrat3):

@saseal

OpenStudy (irishboy123):

you mean \(s(t) = te^t\) ??

OpenStudy (zeesbrat3):

I suppose, I just copied the question honestly. @IrishBoy123

OpenStudy (irishboy123):

if you do mean: \(\large s(t)=te^t\) then \(\large v(t) = \dot s(t)=\frac{d}{dt}( te^t)\) use product rule and set it to zero to find when the things is at rest

OpenStudy (zeesbrat3):

so find the derivative?

OpenStudy (michele_laino):

yes!

OpenStudy (zeesbrat3):

I tried doing what we did but that didn't work, but doing what he just said makes sense. I got 0 as a solution

OpenStudy (michele_laino):

maybe your function is like this: \[\Large s\left( t \right) = t{e^{ - t}}\]

OpenStudy (michele_laino):

sorry, if I compute the first derivative, I got this: \[\Large \frac{{d\left( {t{e^t}} \right)}}{{dt}} = \left( {t + 1} \right){e^t}\]

OpenStudy (michele_laino):

so we have to solve this algebraic equation: \[\Large \left( {t + 1} \right){e^t} = 0\]

OpenStudy (zeesbrat3):

I got -1 as a solution

OpenStudy (michele_laino):

that's right!

OpenStudy (zeesbrat3):

Awesome! Thank you for your help!!

OpenStudy (michele_laino):

:)